OpenStudy (anonymous):

Let f(x) = 2x+3 and g(x) = -x^2 + 5 find ( f of f)(x)

sam (.sam.):

ff(x) f(f(x)) f(2x+3) 2(2x+3)+3

sam (.sam.):

its just like substituting

OpenStudy (anonymous):

Looks like they're giving g(x) to throw you off!

OpenStudy (anonymous):

28 was the rigth answer (:

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

@Lib Libs Does it make sense to plug the eqn of f(x) into f(x)?

OpenStudy (anonymous):

Yes. It does. It's just asking for f of f

OpenStudy (anonymous):

Wait. ; - ; i got confused.

OpenStudy (anonymous):

f(x) = 2x+3) f(f(x)) = f(2x+3) f(2x+3) = 2(2x+3) +3 So you're just plugging your function into itself.

OpenStudy (anonymous):

4x + 3 2x + 3 4x + 9 4x2 + 12x + 9 It would be 2x +3?

OpenStudy (anonymous):

Not quite... f(2x+3) =4x +6 +3 =4x+9

OpenStudy (anonymous):

Darnit D< I will get it right.

OpenStudy (anonymous):

I got two right today :/